What is the mole fraction of the solute that contains 10.0 grams of Phosphorus pentachloride that has been dissolved in 750 mL of water?


1
Expert's answer
2022-02-20T08:40:09-0500

Solution:

solute = phosphorus pentachlorid = PCl5

solvent = water = H2O


The molar mass of PCl5 is 208.24 g/mol

Hence,

Moles of PCl5 = (10.0 g PCl5) × (1 mol PCl5 / 208.24 g PCl5) = 0.048 mol PCl5


The density of water is around approximately 1 g cm−3

Hence,

Mass of H2O = (750 mL H2O) × (1 g H2O / 1 cm3 H2O) × (1 cm3 H2O/ 1 mL H2O) = 750 g H2O

The molar mass of H2O is 18.015 g/mol

Hence,

Moles of H2O = (750 g H2O) × (1 mol H2O / 18.015 g H2O) = 41.632 mol


Mole fraction (χ) of PCl5 = Moles of PCl5 / Total moles of solution

Total moles of solution = Moles of PCl5 + Moles of H2O = 0.048 mol + 41.632 mol = 41.680 mol

Hence,

Mole fraction (χ) of PCl5 = 0.048 mol / 41.680 mol = 0.00115

Mole fraction (χ) of PCl5 = 0.00115


Answer: The mole fraction of the solute (PCl5) is 0.00115